I would like your opinion on the following problem:
A password is composed of 4 different letters (out of 26 upper case letters) and 2 different digits (out of 10 digits from 0 to 9).
How many arrangements are possible?
This my attempt so far.
Attempt 1:
26 x 25 x24 x23 x 10 x 9
Attempt 2:
26C4 x 9C2 x 6

Answer by Theo(13342)

4 different letters and 2 different digits.

if you assumed that the same letter could not be used more than once in the password then you would be correct.

the solution would then be:

26*25*24*23*10*9 which is the same as 26P4*10P2

26P4 is 26*25*24*23
10P2 is 10*9

the combination formula would not be correct since 26C4 = 26*25*24*23 / 1*2*3*4 and 10C2 = 10*9 / 1*2
this is not what you want.

if, however, you allow the same letter to be used in more than one position and the same number to be used in more than one position, then the solution would be:

26*26*26*26*10*10 which could also be written as 26^4*10^2